OpenStudy (anonymous):

Mike is three times as old as his little sister, Mary. In 4 years, he will be twice as old as Mary. How old is Mary?

OpenStudy (anonymous):

Let Mike's age be "3x" yrs. Mary's age is "x" yrs. Mike's age will be "3x+4" yrs. Mary's age will be "x+4" yrs. Mike = 2(Mary) 3x+4 = 2(x+4) 3x+4 = 2x + 16 x = 4 yrs. (Mary's age NOW) 3x = 12 yrs. (Mike's age NOW) Mike is 12 years old.